# Massive Discovery of Low-Dimensional Materials from Universal Computational Strategy

### v1

@@ Abstract
Low-dimensional materials have attractive properties that drive intense efforts for novel materials discovery. However, experiments are tedious for systematic discovery, and present computational methods are often tuned to two-dimensional (2D) materials, overlooking other low-dimensional materials. Here, we combined universal machine-learning interatomic potentials (UMLIPs) and an advanced, interatomic force constant (FC) -based dimensionality classification method to make a massive discovery of novel low-dimensional materials. We first benchmarked UMLIPs’ first-principles-level accuracy in quantifying FCs and calculated phonons for 35,68935{,}689 materials from the Materials Project database. We then used the FC-based method for dimensionality classification to discover $9139$ low-dimensional materials, including $1838$ 0D clusters, $1760$ 1D chains, $3057$ 2D sheets/layers, and $2484$ mixed-dimensionality materials, all of which conventional geometric descriptors have not recognized. By calculating the binding energies for the discovered 2D materials, we also identified $960$ sheets that could be easily or potentially exfoliated from their parent bulk structures.
I Introduction
@@
The discovery of novel functional materials for targeted applications has long been central in materials research. Over the past two decades, the rise of graphene has revealed remarkable properties that have sparked a quest for novel two-dimensional (2D) materials with ever-growing attention Novoselov et al. (2004). To date, the quest has unearthed numerous 2D materials such as transition-metal dichalcogenides (TMDs), MXenes, metallenes, and emerging non‑van der Waals systems, offering promising applications in electronics Radisavljevic et al. (2011); Butler et al. (2013); Manzeli et al. (2017); Xiang et al. (2020); Han et al. (2019), optics Du et al. (2021); Wei et al. (2021); Turunen et al. (2022), magnetism Burch et al. (2018); Gibertini et al. (2019); Klein et al. (2023), sensors Kim et al. (2018); Hosseini-Shokouh et al. (2023); Zhou et al. (2021), catalysis Deng et al. (2016); Cho and Kang (2024); Zhang et al. (2025); Xie et al. (2023), and topological materials Kou et al. (2017); Balandin et al. (2022); Campi et al. (2021); Kezilebieke et al. (2020).
@@
Another approach to materials discovery is computation. Many studies have employed high-throughput computations and data-driven approaches to systematically screen thousands of experimentally known bulk materials, uncovering hundreds to over a thousand exfoliable 2D material candidates Björkman et al. (2012); Lebègue et al. (2013); Ashton et al. (2017); Mounet et al. (2018); Lyngby and Thygesen (2022); Friedrich et al. (2022); Barnowsky et al. (2023). Yet these efforts have been targeted at 2D materials, with only scant studies systematically addressing materials with zero-dimensional (0D), one-dimensional (1D), or mixed-dimensional characteristics Moustafa et al. (2022); Bagheri and Komsa (2023); Li et al. (2024).
@@
To identify materials’ low-dimensional character, most computational methods use geometrical criteria, such as atomic positions and atomic radii Ganose and Jain (2019); Larsen et al. (2019); Cheon et al. (2017); Gorai et al. (2016). This approach has often proven reasonable for dimensionality identification. Geometries are usually calculated by density-functional theory (DFT), which is the workhorse method accurate enough for predicting materials and their properties. Yet even a realistic geometry can be inaccurate for dimensionality prediction, because it relies on empirical correlations between the lengths and strengths of atomic bonds. This shortcoming was recently remedied by a method that uses interatomic force constants (FCs) to identify low-dimensional units from bulk materials Bagheri et al. (2023). While this method—FCDimen—has the merit of relying on chemical bonding quantitatively, so far its applicability has been limited by the high cost of DFT-level calculation of FCs.
@@
Fortunately, universal machine-learning interatomic potentials (UMLIPs) have recently become increasingly helpful in reducing calculation time and generating results at DFT-level accuracy Deringer et al. (2019); Behler (2021); Jinnouchi and Minami (2025). UMLIPS are trained by massive datasets and aim for applicability to all materials that contain nearly all elements. Although trained to predict only energies and forces, UMLIPs have been shown to reliably predict also their derivatives, such as formation energies of bulk materials Riebesell et al. (2025), defect formation energies Berger et al. (2025), and phonon dispersion curves Loew et al. (2025). It would be reasonable to expect that UMLIPs would also be applicable in predicting FCs for fast and reliable identification of low-dimensional units.
@@
Therefore, in this article, we used UMLIPs to calculate FCs and FCDimen to determine the dimensionalities of 35,68935{,}689 materials from the Materials Project (MP) database Jain et al. (2013). To choose the most accurate model, we benchmarked two UMLIPs, MatterSim Yang et al. (2024) and MACE Batatia et al. (2022a, b), against DFT-calculated phonon properties and FCs of an existing phonon database with more than 1010K materials Togo and Tanaka (2015); Loew et al. (2025). We then compared the FCDimen-predicted dimensionalities that DFT and UMLIPs gave for these materials, and concluded that MatterSim is sufficiently accurate for predicting material dimensionalities. Leveraging on MatterSim, we screened 153,234153{,}234 bulk materials and discovered $9139$ low-dimensional materials, including 2D sheets, 1D chains, 0D clusters, and various mixed dimensionalities. Finally, by calculating the interlayer binding energies for the discovered 2D sheets, we found that $960$ of them—all new to the known 2D materials databases—are easily or potentially exfoliable from their parent bulk structure.
II Results
II.1 FCDimen
@@
We begin by briefly reviewing the FC-based dimensionality classification approach FCDimen (see Ref. 38 for details). For a system with potential energy $U(\textbf{r}_{i},...,\textbf{r}_{n})$, where $n$ is the number of atoms and $\textbf{r}_{i}$ the position of atom $i$, the force on atom $i$ is
@@
The actual bonds can then be identified by using the threshold $t={\rm MinFC}$ for the interatomic FCs. In addition, scanning the threshold between $t={\rm MinFC}\ldots{\rm MaxFC}$ gives a score (01)(0-1) for the assigned dimensionality. The FCDimen has proven promising for identifying low-dimensional materials and revealing new non-van der Waals candidates Bagheri et al. (2023).
II.2 Benchmarking
@@
Before using FCDimen to classify dimensionalities, we must first ensure the reliable calculation of FCs. Loew et al. have shown that UMLIPs can predict phonon dispersion curves accurately Loew et al. (2025), with MACE Batatia et al. (2022a, b) and MatterSim Yang et al. (2024) performing the best. We therefore adopted these two UMLIPs for benchmarking [see Supplementary Information (SI) for details]. Reference FCs were taken from the Phonon database Togo and Tanaka (2015), which includes phonon properties at the DFT level for 10,03210{,}032 structures. This database was initially built using the PBEsol exchange-correlation functional. However, because both machine-learning (ML) models were trained on data based on the PBE functional, a fair comparison required the use of Alexandria’s version of the Phonon database that was recalculated using the PBE functional Loew et al. (2025). For clean data and better comparisons, we applied the same screening criteria to the database as Ref. 38. We compared only MaxFC and MinFC, not the entire FC matrix.
@@
MatterSim predicted FCs accurately, with a root mean squared error (RMSE) of $0.64$ eV/Å2 for MaxFC and $0.2$ eV/Å2 for MinFC (Fig. 1a). MACE was slightly less accurate, with RMSEs of $0.86$ eV/Å2 and $0.34$ eV/Å2 (SI Fig. S1a). The FC distributions for ML and DFT are very similar (Fig. 1b and SI Fig. S1b). We remark that the RMSEs between PBE and PBESol were even larger, $0.46$ eV/Å2 for MinFC and $0.99$ eV/Å2 for MaxFC (SI Fig. S2a).
@@
Figure 1: Validation of methodology. Comparing DFT and ML (MatterSim) for: a) MaxFC and MinFC [Equation (5)], b) distribution of MaxFC and MinFC, c) maximum phonon frequencies, and d) number of Phonon database structures in each dimensionality group. Panels (a) and (c) also show the related RMSE.
Figure 2: High-throughput calculation workflow and resulting dataset statistics. (a) Workflow for screening and high-throughput calculations. Each step indicates the number of structures involved. (b) Number of low-dimensional materials discovered in the screened dataset by FCDimen. ”Other” corresponds to mixed dimensionalities with more than two types of dimensionalities (012D, 013D, and 023D). (c) Number of low-dimensional materials classified using the $16$ most common space groups. (d) Heat map with occurrences of each element in low-dimensional materials in the dataset.
In addition to FCs, we also benchmarked FC-derived quantities, including the phonon frequencies and the FCDimen dimensionality predictions. Phonon dispersions and frequencies had already been benchmarked by several UMLIPs, including MatterSim and MACE, which outperformed other UMLIPs Loew et al. (2025). Here, looking at the maximum phonon frequencies, both ML models show excellent agreement with DFT, with RMSE of $0.57$ THz for MatterSim (Fig. 1c) and $0.76$ THz for MACE (SI Fig. S1c).
@@
ML predicted dimensionalities equally reliably. Out of the $4458$ materials screened by the criteria of Ref. 38, MatterSim predicted dimensionalities correctly in $4021$ ($90.2\%$) cases (Fig. 1d). MACE again performed slightly worse, with $3851$ ($86.4\%$) correct dimensionality predictions (SI Fig. S1d). The agreement between PBE- and PBESol-predicted dimensionalities was $90.0\%$ (SI Fig. S2b)—on par with MatterSim accuracy against PBE. Therefore, based on slightly better performance in predicting materials’ FCs and their dimensionalities, we adopt MatterSim as our UMLIP of choice for further calculations.
II.3 High-throughput calculations
@@
Armed with a reliable UMLIP, we may now proceed to high-throughput calculations of FCs to discover new low-dimensional materials in the Materials Project, an approach previously unfeasible due to the high cost of DFT calculations. The calculations were preceded by screening materials in the MP database (Fig. 2a). We excluded materials having more than 100 atoms in the unit cell or including elements unsupported by MatterSim, including noble gases, Tc, Lu, Po, At, and Yb. We also excluded materials containing Lu and La due to unreliable results. To direct our search towards stable and undiscovered low-dimensional materials, we further screened the dataset by the following steps: First, we removed materials already previously found to be low-dimensional, as identified by the method of Larsen et al. Larsen et al. (2019). Second, we removed materials included in the Phonon database since they had already been studied earlier by DFT and FCDimen Bagheri et al. (2023). Third, we removed theoretically predicted materials with energy above the convex hull (Ehull) larger than $0.1$ eV and experimentally observed materials with Ehull larger than $0.2$ eV based on MP data. This screening finally gave us 35,68935{,}689 materials for the UMLIP-assisted high-throughput phonon calculations.
@@
Figure 4: Discovered 2D materials in detail. (a) Comparison of the binding energies of some known 2D materials, calculated using ML (with D3–BJ damping) and DFT (DFT-D3, DF2‑C09, and rVV10). DFT-calculated binding energies using DF2‑C09 and rVV10 are taken from Ref. 27. (b) Number of 2D structures as a function of exfoliation energies using ML D3-BJ. Structures are grouped into easily exfoliable ($E_{b}\leq 35$ meV/Å2), potentially exfoliable ($35<E_{b}<125$ meV/Å2), and strongly bound ($E_{b}>125$ meV/Å2) 2D materials. (c) Correlation between binding energies and MinFC for easily and potentially exfoliable 2D materials. (d) Easily and potentially exfoliable 2D materials categorized into $15$ most common space groups. (e) Chemical structures (top and side views) of selected easily exfoliable 2D materials from the most common species (binary, ternary, and quaternary). Legends indicate space groups, chemical formulas, MP IDs, and binding energies ($E_{b}$ in meV/Å2).
The phonon calculations gave 24,51524{,}515 dynamically stable (no imaginary phonon frequencies >0.1>0.1 THz) materials. Analyzing these materials by FCDimen, we found $9139$ low-dimensional materials including $3057$ sheets (2D), $1760$ chains (1D), $1838$ clusters (0D), and $2484$ mixed dimensionalities (Fig. 2b). Materials with the highest dimensionality scores are presented in SI Tables S1–S4.
The discovered low-dimensional materials could be categorized into $16$ most common space groups, with Pnma, P21/c, and Cmcm being the most common. In addition, binaries, ternaries, and quaternaries were the most commonly appearing species, which can be biased by the nature of the MP database (Fig. 2c). Most materials contain metalloids (Ge and Si), reactive non-metals (O, S, and P), and transition metals (Fe, Co, Ni) (Fig. 2d). Promethium has the least number of compounds, and oxides dominate the materials, possibly because most of the materials in MP are oxides.
[…]
@@
To investigate the discovered 2D materials in more detail, we calculated their interlayer binding energies $E_{b}$. We accounted for the weak van der Waals interactions by applying the D3 dispersion correction Grimme (2006); Grimme et al. (2010, 2011); Takamoto et al. (2022). We ensured the reliability of ML-D3 calculations by benchmarking our results against DFT-calculated binding energies for a group of known 2D materials (Fig. 4a). The binding energies from ML (D3-BJ damping), DFT (D3-BJ, DF2-C09 Mounet et al. (2018), and rVV10 Mounet et al. (2018)) agree reasonably well. However, most of the D3-BJ binding energies (ML and DFT) are above the two other approaches due to the functional form or damping. We therefore selected our exfoliation energy thresholds slightly higher ($\sim 5$ meV/Å2) than those in Refs. 24; 27; 54.
@@
We grouped the discovered $3057$ 2D materials into three categories based on $E_{b}$: easily exfoliable ($E_{b}\leq 35$ meV/Å2), potentially exfoliable ($35<E_{b}<125$ meV/Å2), and strongly bound ($E_{b}>125$ meV/Å2) (Fig. 4b). Consequently, we found $183$ easily exfoliable and $953$ potentially exfoliable 2D materials. Apart from BN (mp-2653), all the easily exfoliable materials had MinFC below $5$ eV/Å2 (Fig. 4c).
@@
The 2D materials are categorized into the $15$ most common space groups, with Cmcm, P21/c, and C2/m being the most prominent (Fig. 4d). Fig. 4e shows the chemical structures of six common species as representative prototypes. Phonon dispersions of these representatives are shown in SI Fig. S3. We compared the chemical structures and formulas of the discovered 2D materials in known databases, including computational two-dimensional database (C2DB) Haastrup et al. (2018); Gjerding et al. (2021), materials cloud two-dimensional structure database (MC2D) Mounet et al. (2018); Campi et al. (2023), 2DMatPedia Zhou et al. (2019), 2D topological insulators Marrazzo et al. (2019), NOMAD (RAE database) Evans et al. (2024), and Materials Project Jain et al. (2013); Ganose and Jain (2019). We found $152$ easily exfoliable (SI Table S7) and $808$ potentially exfoliable 2D materials that could not be found in these databases.
III Discussion

### v2

@@ Abstract
Low-dimensional materials have attractive properties that drive intense efforts for novel materials discovery. However, experiments are tedious for systematic discovery, and present computational methods are often tuned to two-dimensional (2D) materials, overlooking other low-dimensional materials. Here, we combined universal machine-learning interatomic potentials (UMLIPs) and an advanced, interatomic force constant (FC) -based dimensionality classification method to make a massive discovery of novel low-dimensional materials. We first benchmarked UMLIPs’ first-principles-level accuracy in quantifying FCs and calculated phonons for 35,68935{,}689 materials from the Materials Project database. We then used the FC-based method for dimensionality classification to discover $9139$ low-dimensional materials, including $1838$ 0D clusters, $1760$ 1D chains, $3057$ 2D sheets/layers, and $2484$ mixed-dimensionality materials, all of which conventional geometric descriptors have not recognized. By calculating the binding energies for the discovered 2D materials, we also identified $887$ sheets that could be easily or potentially exfoliated from their parent bulk structures.
Introduction
@@
The discovery of novel functional materials for targeted applications has long been central in materials research. Over the past two decades, the rise of graphene has revealed remarkable properties that have sparked a quest for novel two-dimensional (2D) materials with ever-growing attention [Novoselov2004]. To date, the quest has unearthed numerous 2D materials such as transition-metal dichalcogenides (TMDs), MXenes, metallenes, and emerging non‑van der Waals systems, offering promising applications in electronics [Radisavljevic2011, Butler2013, Manzeli2017, Xiang20_Sci, Han19_NComm], optics [Du21_NComm, Wei21_AM, Turunen22_NRevPhys], magnetism [Burch2018, Gibertini2019, Klein23_ACSNano], sensors [Kim_2018, mx_sens, Zhou_2021], catalysis [Deng2016, D4NR00147H, Zhang_2025, Xie2023], and topological materials [Kou2017, Balandin22_MT, Campi2021, Kezilebieke2020].
@@
Another approach to materials discovery is computation. Many studies have employed high-throughput computations and data-driven approaches to systematically screen thousands of experimentally known bulk materials, uncovering hundreds to over a thousand exfoliable 2D material candidates [Bjorkman2012, PRX2013, Ashton2017, Mounet2018, Lyngby2022, Arkady2022, Barnowsky2023, barnowsky2025l]. Yet these efforts have been targeted at 2D materials, with only scant studies systematically addressing materials with zero-dimensional (0D), one-dimensional (1D), or mixed-dimensional characteristics [Hadeel2022, Screening2022, Li2024, Khazaei_lowD].
@@
To identify materials’ low-dimensional character, most computational methods use geometrical criteria, such as atomic positions and atomic radii [MP_dimen, Larsen_2019, Cheon2017, Gorai2016]. This approach has often proven reasonable for dimensionality identification. Geometries are usually calculated by density-functional theory (DFT), which is the workhorse method accurate enough for predicting materials and their properties. Yet even a realistic geometry can be inaccurate for dimensionality prediction, because it relies on empirical correlations between the lengths and strengths of atomic bonds. This shortcoming was recently remedied by a method that uses interatomic force constants (FCs) to identify low-dimensional units from bulk materials [FCDimen]. While this method—FCDimen—has the merit of relying on chemical bonding quantitatively, so far its applicability has been limited by the high cost of DFT-level calculation of FCs.
@@
Fortunately, universal machine-learning interatomic potentials (UMLIPs) have recently become increasingly helpful in reducing calculation time and generating results at DFT-level accuracy [Deringer19_AM, Behler2021_review, Jinnouchi2025_review]. UMLIPS are trained by massive datasets and aim for applicability to all materials that contain nearly all elements. Although trained to predict only energies and forces, UMLIPs have been shown to reliably predict also their derivatives, such as formation energies of bulk materials [Riebesell2024matbench], defect formation energies [Burger2025], and phonon dispersion curves [alexandria]. It would be reasonable to expect that UMLIPs would also be applicable in predicting FCs for fast and reliable identification of low-dimensional units.
@@
Therefore, in this article, we used UMLIPs to calculate FCs and FCDimen to determine the dimensionalities of 35,68935{,}689 materials from the Materials Project (MP) database [mp]. To choose the most accurate model, we benchmarked two UMLIPs, MatterSim [MatterSim] and MACE [mace1, mace2], against DFT-calculated phonon properties and FCs of an existing phonon database with more than 1010K materials [Togo2015, alexandria]. We then compared the FCDimen-predicted dimensionalities that DFT and UMLIPs gave for these materials, and concluded that MatterSim is sufficiently accurate for predicting material dimensionalities. Leveraging on MatterSim, we screened 153,234153{,}234 bulk materials and discovered $9139$ low-dimensional materials, including 2D sheets, 1D chains, 0D clusters, and various mixed dimensionalities. Finally, by calculating the interlayer binding energies for the discovered 2D sheets, we found that $887$ of them—all new to the known 2D materials databases—are easily or potentially exfoliable from their parent bulk structure.
Methods
Computational methods
The machine-learning (ML) calculations were performed using MatterSim with pre-trained model v1.0.0-5M [MatterSim] and MACE [mace1, mace2] with the pre-trained large model of MACE-MP-0 (MACE-MP-0b3) without dispersion. The MatterSim-v1 model is based on the M3GNet [m3gnet] architecture, which was trained on relaxations obtained from the Materials Project [mp]. Both UMLIP models are available as calculators in the Atomic Simulation Environment (ASE) [ase-paper]. ASE was used, with the help of Python, to construct the calculation workflow. In the benchmark, all parameters were kept the same as those in the Phonon database [Togo2015]. All materials were relaxed using the FIRE [FIRE] and BFGS algorithm to forces below $1$ meV/Å. The crystal structures and their properties were extracted using MP-API [MP_API] and Python materials genomics (Pymatgen) [pymatgen]. The space groups are calculated using spglib [spglib] with symmetry precision $10^{-6}$.
Initial dimensionalities were calculated using Larsen [Larsen_2019] method, which is implemented in ASE [ase-paper]. Phonon calculations were done using the Phonopy package [phonopy1, phonopy2], using a displacement of $0.01$ Å. The supercell sizes used for determining the force constants were chosen so that the number of atoms was less than or equal to $250$ while keeping the point groups of the Bravais lattices of the conventional unit cell and the supercell the same.
Initial structures of known 2D materials (Figure 4a) were extracted from Ref. [Mounet2018]. All density-functional theory calculations were performed using VASP (Vienna Ab initio Simulation Package) software [Kresse1996, PhysRevB.54.11169], together with the projector-augmented wave method [PhysRevB.50.17953]. The Becke-Johnson (BJ) damping variant of the DFT-D3 functional was adopted [Grimme_BJ]. The plane wave cutoff is set to $520$ eV. The Brillouin zone of the unit cell is sampled by a $\Gamma$-centered k-point mesh whose density is defined by $R_{k}=20$, which in VASP determines the subdivisions $N_{1}$, $N_{2}$, and $N_{3}$ along the reciprocal lattice vectors $b_{1}$, $b_{2}$, and $b_{3}$, respectively, via $N_{i}={\rm max}(1,R_{k}|b_{i}|+0.5)$ and rounded to an integer. The energy convergence criterion between consequential steps is set to $10^{-8}$ eV for the self-consistent loop and $10^{-6}$ eV for ionic relaxation. Structures were considered to be relaxed when no force exceeded $1$ meV/Å.
In the machine-learning binding energy calculation, we employed torch-dftd3 [Grimme2006, Grimme2010, Grimme_BJ, Takamoto2022] alongside MatterSim to add the D3 correction in our calculations to correct the weak van der Waals (vdW) interaction. This approach has already proven successful for the investigation of 2D material heterobilayers [Kristian2025].
Binding energy calculation
The method used to obtain the binding energy $E_{b}$ is adapted from Ref. [Bjorkman2012]. Taking materials from MP, the bulk unit cell is first relaxed with D3 correction to obtain the initial energy $E_{0}$ as well as the force constant. Using the latter, 2D units are identified using FCDimen. The resulting monolayers are then isolated and relaxed, leading to the layer energies $E_{l}$. The binding energy $E_{b}$ is then found as
$E_{b}=\frac{\sum E_{l}-E_{0}}{A\cdot N_{l}},$ (1)
Results
FCDimen
@@ Results
We begin by briefly reviewing the FC-based dimensionality classification approach FCDimen previously introduced in Ref. [FCDimen]. For a system with potential energy $U(\textbf{r}_{i},...,\textbf{r}_{n})$, where $n$ is the number of atoms and $\textbf{r}_{i}$ the position of atom $i$, the force on atom $i$ is
@@
To identify low-dimensional units, we treat the atoms as nodes in a graph. A pair of atoms is considered connected when the interatomic FCs are larger than a threshold $t={\rm MinFC}$, i.e., when $\Phi_{ij}\geq t$. The low-dimensional units are then defined as the resulting connected components of atoms linked by bonds. The dimensionality of these components is determined by the periodic extension of their unit cell (e.g., a component connected to its own periodic image in two directions is classified as 2D). Furthermore, scanning the threshold between $t={\rm MinFC}\ldots{\rm MaxFC}$ gives a score in the range 0–1 for the assigned dimensionality. The FCDimen has proven promising for identifying low-dimensional materials and revealing new non-van der Waals candidates (see Ref. [FCDimen] for more details).
Benchmarking
@@
Before using FCDimen to classify dimensionalities, we must first ensure the reliable calculation of FCs. Loew et al. have shown that UMLIPs can predict phonon dispersion curves accurately [alexandria], with MACE [mace1, mace2] and MatterSim [MatterSim] performing the best. We therefore adopted these two UMLIPs for benchmarking [see Methods for details]. Reference FCs were taken from the Phonon database [Togo2015], which includes phonon properties at the DFT level for 10,03210{,}032 structures. This database was initially built using the PBEsol exchange-correlation functional. However, because both machine-learning (ML) models were trained on data based on the PBE functional, a fair comparison required the use of Alexandria’s version of the Phonon database that was recalculated using the PBE functional [alexandria]. For clean data and better comparisons, we applied the same screening criteria to the database as Ref. [FCDimen]. Screening involved the following criteria: (i) enforcing a minimum supercell lattice constant of 5 Å; (ii) requiring that FCs at half the lattice distance remain below 20% of the MaxFC to ensure the supercell size was sufficient for force constant (FC) decay; and (iii) filtering out chemically inert noble gases with negligible forces. Finally, we restricted our dataset to dynamically stable structures by excluding any materials exhibiting imaginary phonon frequencies (no imaginary phonon frequencies >0.1>0.1 THz) as determined by either DFT or ML. We compared only MaxFC and MinFC, not the entire FC matrix.
@@
MatterSim predicted FCs accurately, with a root mean squared error (RMSE) of $0.64$ eV/Å2 for MaxFC and $0.2$ eV/Å2 for MinFC (Figure 1a). MACE was slightly less accurate, with RMSEs of $0.86$ eV/Å2 and $0.34$ eV/Å2 (Supplementary Information (SI) Figure S1a). The FC distributions for ML and DFT are very similar (Figure 1b and SI Figure S1b). We remark that the RMSEs between PBE and PBESol were even larger, $0.46$ eV/Å2 for MinFC and $0.99$ eV/Å2 for MaxFC (SI Figure S2a). Note that these correspond to relative errors of 5.55.5% for MaxFC and 9.29.2% for MinFC, which should ensure accurate predictions of the dimensionalities when combining UMLIPs with FCDimen.
@@
Figure 1: Validation of methodology. Comparing DFT and ML (MatterSim) for: a) MaxFC and MinFC [Equation (6)], b) distribution of MaxFC and MinFC, c) maximum phonon frequencies at $\Gamma$-point, and d) number of Phonon database structures in each dimensionality group. Panels (a) and (c) also show the related RMSE.
@@
In addition to FCs, we also benchmarked FC-derived quantities, including the phonon frequencies and the FCDimen dimensionality predictions. Phonon dispersions and frequencies had already been benchmarked by several UMLIPs, including MatterSim and MACE, which outperformed other UMLIPs [alexandria]. Here, looking at the maximum phonon frequencies at $\Gamma$-point, both ML models show excellent agreement with DFT, with RMSE of $0.57$ THz for MatterSim (Figure 1c) and $0.76$ THz for MACE (SI Figure S1c). ML predicted dimensionalities equally reliably. Out of the $4458$ materials screened by the criteria of Ref. [FCDimen], MatterSim predicted dimensionalities correctly in $4021$ ($90.2\%$) cases (Figure 1d). MACE again performed slightly worse, with $3851$ ($86.4\%$) correct dimensionality predictions (SI Figure S1d). The agreement between PBE- and PBESol-predicted dimensionalities was $90.0\%$ (SI Figure S2b)—on par with MatterSim accuracy against PBE. Therefore, based on slightly better performance in predicting materials’ FCs and their dimensionalities, we adopt MatterSim as our UMLIP of choice for further calculations.
High-throughput calculations
@@
Armed with a reliable UMLIP, we may now proceed to high-throughput calculations of FCs to discover low-dimensional materials in the Materials Project, an approach previously unfeasible due to the high cost of DFT calculations. The calculations were preceded by screening materials in the MP database (Figure 2a). We excluded materials having more than 100 atoms in the unit cell or including elements unsupported by MatterSim, including noble gases, Tc, Lu, Po, At, and Yb. We also excluded materials containing Lu and La due to unreliable results. To direct our search towards stable and undiscovered low-dimensional materials, we further screened the dataset by the following steps: First, we removed materials already previously found to be low-dimensional, as identified by the method of Larsen et al. [Larsen_2019], as demonstrated in our previous work (Ref. [FCDimen]), the materials identified as low-dimensional by conventional geometric descriptors are also found to be low-dimensional by FCDimen. Second, we removed materials included in the Phonon database since they had already been studied earlier by DFT and FCDimen [FCDimen]. Third, we removed theoretically predicted materials with energy above the convex hull (Ehull) larger than $0.1$ eV and experimentally observed materials with Ehull larger than $0.2$ eV based on MP data. This screening finally gave us 35,68935{,}689 materials for the UMLIP-assisted high-throughput phonon calculations.
Figure 2: High-throughput calculation workflow and resulting dataset statistics. (a) Workflow for screening and high-throughput calculations. Each step indicates the number of structures involved. (b) Number of low-dimensional materials discovered in the screened dataset by FCDimen. "Other" corresponds to mixed dimensionalities with more than two types of dimensionalities (012D, 013D, and 023D). (c) Number of low-dimensional materials classified using the $16$ most common space groups. (d) Heat map with occurrences of each element in low-dimensional materials in the dataset.
The phonon calculations gave 24,51524{,}515 dynamically stable (no imaginary phonon frequencies >0.1>0.1 THz) materials. Analyzing these materials by FCDimen, we found $9139$ low-dimensional materials including $3057$ sheets (2D), $1760$ chains (1D), $1838$ clusters (0D), and $2484$ mixed dimensionalities (Figure 2b). Materials with the highest dimensionality scores are presented in SI Tables S1–S4.
The discovered low-dimensional materials could be categorized into $16$ most common space groups, with Pnma, P21/c, and Cmcm being the most common. In addition, binaries, ternaries, and quaternaries were the most commonly appearing species, which can be biased by the nature of the MP database (Figure 2c). Most materials contain metalloids (Ge and Si), reactive non-metals (O, S, and P), and transition metals (Fe, Co, Ni) (Figure 2d). Promethium has the least number of compounds, and oxides dominate the materials, possibly because most of the materials in MP are oxides.
[…]
@@
To investigate the discovered 2D materials in more detail, we calculated their interlayer binding energies $E_{b}$. We accounted for the weak van der Waals interactions by applying the D3 dispersion correction [Grimme2006, Grimme2010, Grimme_BJ, Takamoto2022]. Recently, Sauer et al.  [Kristian2025] demonstrated that UMLIPs with dispersion corrections are ready for large-scale structural simulations of 2D vdW material heterostructures, identifying MatterSim as one of the leading models in their evaluation. Consistent with these findings, we validated the reliability of ML-D3 calculations by benchmarking our results against DFT-calculated binding energies for a group of known 2D materials (Figure 4a). The binding energies from ML (D3-BJ damping), DFT (D3-BJ, DF2-C09 [Mounet2018], and rVV10 [Mounet2018]) agree reasonably well. However, most of the D3-BJ binding energies (ML and DFT) are above the two other approaches due to the functional form or damping. We therefore selected our exfoliation energy thresholds slightly higher ($\sim 5$ meV/Å2) than those in Refs. [Bjorkman2012, Mounet2018, Marzari2023].
Figure 4: Discovered 2D materials in detail. (a) Comparison of the binding energies of some known 2D materials, calculated using ML (with D3–BJ damping) and DFT (DFT-D3, DF2‑C09, and rVV10). DFT-calculated binding energies using DF2‑C09 and rVV10 are taken from Ref. [Mounet2018]. (b) Number of 2D structures as a function of exfoliation energies using ML D3-BJ. Structures are grouped into easily exfoliable ($E_{b}\leq 35$ meV/Å2), potentially exfoliable ($35<E_{b}<125$ meV/Å2), and strongly bound ($E_{b}>125$ meV/Å2) 2D materials. (c) Correlation between binding energies and MinFC for easily and potentially exfoliable 2D materials. (d) Easily and potentially exfoliable 2D materials categorized into $15$ most common space groups. (e) Chemical structures (top and side views) of selected easily exfoliable 2D materials from the most common species (binary, ternary, and quaternary). Legends indicate space groups, chemical formulas, MP IDs, and binding energies ($E_{b}$ in meV/Å2).
@@
We grouped the discovered $3057$ 2D materials into three categories based on $E_{b}$: easily exfoliable ($E_{b}\leq 35$ meV/Å2), potentially exfoliable ($35<E_{b}<125$ meV/Å2), and strongly bound ($E_{b}>125$ meV/Å2) (Figure 4b). Consequently, we found $183$ easily exfoliable and $953$ potentially exfoliable 2D materials. Apart from BN (mp-2653), all the easily exfoliable materials had MinFC below $5$ eV/Å2 (Figure 4c).
@@
The 2D materials are categorized into the $15$ most common space groups, with Cmcm, P21/c, and C2/m being the most prominent (Figure 4d). Figure 4e shows the chemical structures of the six selected 2D materials from the most common species (binary, ternary, and quaternary). We performed DFT phonon calculations for four non-magnetic 2D materials and compared them with ML-predicted phonon dispersions (SI Figure S3). Although the agreement for these particular cases is not strong, it validated the accuracy of UMLIPs in identifying true stability, consistent with Ref. [alexandria]. ML predicted phonon dispersions of these selected 2D materials are shown in SI Figure S4. We compared the chemical structures, formulas, and Materials IDs (if applicable) of the discovered 2D materials in known databases, including computational two-dimensional database (C2DB) [c2db1, c2db2], materials cloud two-dimensional structure database (MC2D) [Mounet2018, Marzari2023], 2DMatPedia [Zhou2019], 2D topological insulators [topo], NOMAD (RAE database) [nomad1], 2D charged building block databases (2DBBs) [Pan2023], and Materials Project [mp, MP_dimen]. We found $146$ easily exfoliable (SI Table S7) and $741$ potentially exfoliable 2D materials that could not be found in these databases.
Conclusion